OpenStudy (anonymous):

Algebra help!!! A group of 19 students want to see the show at the planetarium. Tickets cost $11 for each student who is a member and $13 for those who are not. Joining the program to become a member costs $5 a year. Write and equation that can be used to determine n, the number of visits per year for which cost of being a member is equal to the cost of not being a member.

OpenStudy (madhu.mukherjee.946):

total no. of student =19 let members =n non- members=19-n so members would pay=11n non-members would pay 13(19-n) and if a non member wants to become a member then he has to$5 ,so in total he will pay 11n+5 so the equation is- 11n+(11n+5)=13(19-n)

OpenStudy (anonymous):

Oh my gosh thank you so much!!! Been trying to get this equation right for the past hour. I kept messing up and not putting 11n+5

OpenStudy (madhu.mukherjee.946):

your welcome